OpenStudy (clamin):

If F(x)= x-2, which of the following is the inverse of F(x)? a. F^-1(x)= 2-x b. F^-1(x) = x-2 c. F^-1(x)= x+2 d. F^-1(x)=2x

OpenStudy (greatlife44):

f(x) = x-2 y = x-2 x = y-2 x+2 = y = inverse